OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1

OpenStudy (anonymous):

Try solving through substitution. −2x + y = 1 is the same as y = 2x + 1 -4x + 2x + 1 = -1 -2x = -2 x = 1 Now plug 1 in for x and solve for y.

OpenStudy (anonymous):

To solve this, there are a number of methods. The easiest way here is to solve with elimination. Multiply one of the equations by -1 and then the y's can cancel out. -2x+y=1 -4x+y=-1 Turns into 2x-y=-1 -4x+y = -1 Now you can just add the two equations together to find x, and then put the value of x back into one of the equations to find y. Once you have that, the solution is written as an ordered pair: (x,y)